Problem 4:(15 points) Justin Clement Company has had the following pattern earnings per share over the last five years: Earnings per Share Year 2014 4.00 20154.20 2016 2017 2018 4.41 4.63 4.86 earnings per share have grown at a constant rate (on a rounded basis) and will continue to d future. The payout ratio is 40%. Estimate the earnings per share at the end of 2019 Estimate dividends per share at the end of 2019 If the required rate of return is 13%, what is the estimated value of the stock price Po at end of 2018?
Answer #1

A) Earning Per Share (EPS) is increasing at a constant rate of 5%. So EPS for 2019 would be euro 4.86 * 1.05 = Euro 5.103

Calculation to show that EPS is increasing at constant rate of 5% is as under

Euro (4.20 - 4.0) / 4.0 = 5%

(4.41 - 4.20)/4.20 = 5%

B) Payout ratio is 40% that means 40% of earning per share is paid out as dividend to investors. Therefore Dividend per share for 2019 is

5.103 * 0.4 = Euro 2.04

Formula of Payout ratio = [Dividend Paid per share / Earning per share] * 100

C) Stock price = D1/ (r-g)

Where D1 is Dividend per share at t1 in this case 2019 since t is 2018

r - required rate of return given as 13%

g - growth rate in Dividends. Here this is same as growth rate in EPS. i.e 5%

Therefore Stock price = 2.04 / (13%-5%) = 2.04/0.08

Stock price = Euro 25.5 at the end of 2018